OpenStudy (anonymous):

Is this an example of permutation or combination? In how many different ways could a committee of 5 people be chosen from a class of 30 students?

OpenStudy (anonymous):

KK i'm not completely sure if this is how you do it but ill give it a shot.

OpenStudy (anonymous):

You can multiply 30x29x28x27x26x25x24........... all the way to 1. There should be a button on the calculator that will make this step shorter. You do the same with 5x4x3x2x1. When you get your two answers you'll divide them together, and this will give you the amount of combinations. I think....

OpenStudy (anonymous):

If you're just CHOOSING things, and it ends there, then it's a combination. If you're going to ARRANGE them in some way, then it's a permutation.

OpenStudy (anonymous):

@Brent0423 did this help at all?

OpenStudy (anonymous):

Mr. Man lol It won't, because your solution gives far too large a number :)

OpenStudy (anonymous):

Actually, Mr. Man, your "solution" gives the number of ways you can arrange twenty-five people in a row, from a choice of thirty students. That number is like... really huge :D

OpenStudy (anonymous):

so its a combination?

OpenStudy (anonymous):

is that correct @PeterPan

OpenStudy (anonymous):

It is a combination, yes :)

OpenStudy (anonymous):

But when you divide it, wont it give you a small enough answer.

OpenStudy (anonymous):

Try it... just for lawlz :D \[\frac{30\cdot29\cdot28...\cdot 6\cdot5\cdot4\cdot3\cdot2\cdot 1}{5\cdot4\cdot3\cdot 2\cdot1}\] \[\frac{30\cdot29\cdot28...\cdot 6\cdot\cancel{5\cdot4\cdot3\cdot2\cdot 1}}{\cancel{5\cdot4\cdot3\cdot 2\cdot1}}\] \[\large 30\cdot29\cdot28\cdot...7\cdot6\] I don't know about you, but that number's pretty darn huge

OpenStudy (anonymous):

yea i guess haha.